OpenStudy (anonymous):

Marty has a piggy bank that contains $2.25. The piggy bank contains only quarters and dimes. There are twice as many dimes as quarters in the bank.How many dimes are in the piggy bank?

OpenStudy (anonymous):

10d + 25q = 225 d = 2q 10(2q) +25q = 225 20q + 25q = 225 45q = 225 q = 5

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

i need to know how many dimes there are